OpenStudy (anonymous):

Can someone explain implicit differentiation to me using my example problem?

OpenStudy (anonymous):

\[x ^{3}-xy+y ^{2}=7\]

OpenStudy (tkhunny):

If y = f(x), then y' = f'(x). Not much more to it that that, really. Learn your Chain Rule THOROUGHLY, first!

OpenStudy (abmon98):

x^3-xy+y^2=7 take derivative of both sides in respect of x x^3 differentiate with respect to x which gives us 3x^2 -xy thats differentiated using product rule. first factor * the derivative of the second factor) x*y'+y*x') Derivative of y (chain rule) 2yy' differentiation of 7 is 0 3x^2-xy'-yx'+2yy'=0 3x^2-yx'=xy'-2yy' 3x^2-yx'=y'(x-2y) 3x^2-yx'/x-2y=y'

OpenStudy (anonymous):

Wow. That actually makes a lot of sense. Much more simple than I thought it would be too. Thanks!

OpenStudy (abmon98):

good to hear that, your most welcome :D